# OCR Hinders RAG: Evaluating the Cascading Impact of OCR on Retrieval-Augmented Generation

### v2

@@ 4.3 In-depth analysis of OCR noise’s impact on RAG
In this section, we conduct an in-depth analysis of the impact of Semantic Noise and Formatting Noise on RAG systems, using perturbed structured data with varying levels of perturbations. For each type of OCR noise, we introduce three noise levels—mild, moderate, and severe—to systematically assess their effects. Since our findings in Sec. 4.2 indicate that edit distance fails to accurately capture the degree of OCR noise, we instead define the ratio $r_{\text{noise}}$, representing the proportion of Q&A pairs affected by OCR noise as a measure. Specifically, for each Q&A pair, we compute the LCS between its evidence context and the corresponding perturbed structured data. If LCS exceeds 0.95, the Q&A pair is considered as unaffected, otherwise, it is affected. We then use the ratio ($r_{\text{noise}}$) to quantify the degree of perturbations, with 0 representing to ground truth structured data and values approaching 1 indicating greater perturbation. This approach allows us to align the perturbation levels of Formatting Noise and Semantic noise for fair comparisons. Additionally, for Formatting Noise, we evaluate its retrieval performance with modified LCS calculations by excluding stylistic commanding introduced during perturbation, ensuring a fair assessment of retrieval accuracy. For each degree of Semantic Noise, we report the average RAG performance using three different OCR results, including MinerU, GOT, and Qwen2.5-VL-72B.
@@ 4.3.1 Fine-grained impact on retrieval and generation
Semantic Noise significantly influences both retrieval and generation phases. As illustrated in Fig. 3, increasing Semantic Noise from mild ($r_{\text{noise}}=0$) to severe ($r_{\text{noise}}>0.6$) results in nearly a $50\%$ performance decline for most retrievers and LLMs. In the retrieval stage, both the sparse retriever BM25 and the dense retriever BGE-M3 suffer consistent performance declines across all types of questions, suggesting that dense retrieval’s stronger comprehension does not provide robustness against Semantic Noise. In the generation phase, all LLMs struggle with Semantic Noise, among which performance on reading-order decreases the most. Interestingly, although the way we introduce Semantic Noise should primarily affect text recognition, questions related to multimodal elements (tables, formulas and charts) degrade even further, highlighting the challenges in parsing, understanding and reasoning over multimodal document data.
@@ 4.3.2 Impact on end-to-end evaluation
Semantic Noise consistently demonstrate a strong impact, while Formatting Noise affects specific retrievers and LLMs differently. Semantic Noise consistently degrades performance across all combination of retrievers and LLMs, particularly on multimodal questions involving tables, formulas, and charts. In contrast, the effect of Formatting Noise is more variable when using smaller LLMs, Llama3.1-8B and Qwen2-7B, for generation, despite greatly reduced retrieval accuracy, the overall performance shows a slight change due to their limited information integration capabilities. Conversely, using a larger LLM, Qwen2-72B, is highly sensitive to retrieval performance. While its generation performance shows light changes, the overall performance decreases more, especially on questions related to multimodal elements.
In summary, Semantic Noise significantly affects each stage of RAG and the entire system. The impact of Formatting Noise varies with different retrievers and language models, particularly affecting questions related to multimodal elements.
@@ 4.3.3 Impact of table format
In addition to perturbations, we investigate the influence of different table formats as a kind of Formatting Noise. As shown in Fig. 4, HTML tables show inferior performance during retrieval compared to the Markdown and LaTeX formats. Markdown and LaTeX formats perform similarly, with BGE-M3 demonstrating a better understanding of Markdown. In the generation phase, HTML and LaTeX showed similar performance across all models, but the Markdown format performed worse due to the lack of support for merging cells. In end-to-end evaluations affected by low retrieval performance, using HTML tables is comparatively worse, while the combination of Qwen2-72B and BGE-M3 using Markdown achieves the best performance.
@@ 4.3.4 Error analysis
We further conduct error analysis to understand the bottlenecks of the RAG system with OCR noise in a quality approach. Specifically, we calculate the distribution of OCR errors and RAG errors when using ground truth structured data and perturbed data with severe Formatting Noise and Semantic Noise. Our evaluation employs BGE-M3 as the retriever and assesses error distributions using two generators: Qwen2-7B and Qwen2-72B. We use the same strategies to identify the proportion of OCR errors as in Sec. 4.3. The distribution of these errors is illustrated in Fig. 5. It indicates that when the proportion of OCR errors is similar for Semantic Noise and Formatting Noise (66% vs. 61%), it performs better with Formatting Noise. Of these, about half of the correct responses when using Qwen2-7B as a generator were done by the model with incorrect OCR results. Meanwhile, larger LLMs are more robust to OCR noise. Compared to Qwen2-7B, Qwen-72B has a 9.4% and 3.5% higher percentage of samples with OCR errors but ultimately correct in Formatting Noise and Semantic Noise, respectively.
@@ I Instruction Prompts
Q&A Generation Prompt Template. The template is shown in Tab. S6. Following [44, 49], we instruct GPT-4o to generate questions with clear entities and require three levels of difficulty for question diversity.
RAG Generation Stage Prompt Template. The prompt template for LLMs and VLMs with text-only input is shown in Tab. S9.
Vision-Language Models OCR Prompt Template. We tune the prompt for the best performance of VLMs OCR, by comparing simple and detailed instructions as shown in Tab. S10. Results in Tab. S11 indicate that the detailed prompt consistently performs better across all evaluations, so it is used by default.
@@ II.1 Document details
DUDE: We extract documents from the validation and test splits of DUDE, applying manual screening based on our criteria Fig. 1 to exclude samples infeasible for structured data parsing and classify each of them into 7 domains. We finally selected 450 PDFs with 4,058 images from 2,069 PDF candidates.
@@
Each document is manually reviewed by primary authors to ensure its availability for academic use. Detailed domain statistics are shown in Tab. S1
@@ II.4 Q&A generation
To generate high-quality Q&A pairs covering diverse tasks and evidence sources, we define multiple prompts for each task, as detailed Tabs. S6, S7 and S8. For Chinese questions, we provide the same set of templates in Chinese to ensure that the model generates Chinese responses. Q&A with different evidence sources. For Q&A generation with evidence sourced from plaint text, table, formula and chart, we extract relevant pages from the ground truth structured data and use GPT-4o to generate Q&A pairs grounded in the corresponding evidence via tailored prompts. For Q&A related to reading order, we leverage MinerU [41], the leading model for reading order recognition [31], to identify the reading order and bounding box of paragraphs in each document. When working with documents from OmniDocBench [31], we directly use the ground truth reading order from its annotations. We verify the layout detection and reading order predictions, selecting paragraph pairs that meet one of the following criteria:
@@
We then randomly sample 1,500 candidate matches, manually correcting approximately 20% where MinerU’s predictions are inaccurate. We then prompt GPT-4o to generate Q&A pairs using the prompts in Tab. S7. We find that this simple prompting-based strategy can effectively generate questions with diverse evidence sources, with over 90% correctly aligned with their evidence source in our Q&A verification process.
Q&A with different tasks. To generate both understanding and reasoning questions, we apply the corresponding prompts from Tab. S7. For multi-page Q&A generation, we employ two different approaches to generate Q&A candidates: (1) Combine questions from two single-page Q&As that mention the same entity. (2) Generating multi-page questions from two paragraphs on different pages that reference the same entity. Specifically, we use spaCy [17] for named entity recognition in both single-page Q&As and document paragraphs. We then filter out candidate pairs, including: (1) Single-page Q&A pairs where the entity in one answer appears in another question. (2) Paragraph pairs that share the same entity. We finally utilize the prompts in Tab. S8 to generate multi-page questions. However, despite the many optimizations of the prompt and generation strategies we tried, GPT-4o sometimes produces Q&A pairs that are either answerable with a single paragraph or simply concatenate two single-page questions while maintaining separate evidence sources instead of high-quality and realistic multi-page Q&As. To address these limitations, we develop a comprehensive filtering process to ensure the quality of multi-page Q&As, as detailed in Sec. II.5.
@@
Faithfulness to Task Definition. Based on the Q&A verification prompts in [10], we use the prompts in Tab. S12 to assess faithfulness using DeepSeek-V3. To verify the validity of evidence sources, we locate them in the original ground truth structured data and ensure they originate from the correct corresponding LaTeX code environments. For the multi-page and reading-order questions, we employ GPT-4o to generate three responses: (1) without context, (2) with context A, and (3) with context B. If any response yields a correct answer, the question is excluded, ensuring that only truly multi-page or reading-order-related questions remain.
@@ III.2 Rules for Semantic Noise introduction
  • Random Rotation: Apply a random rotation of $-3^{\circ}$ and $+3^{\circ}$.
@@
We generate three document image datasets with varying noise levels and parse structured data using MinerU, GOT, and Qwen2.5-VL, resulting in 9 perturbed datasets. The examples of distorted documents are shown in Fig. S1. In Sec. 4.3, we evaluate RAG performance on these datasets, reporting the average results for each noise level.
@@ IV.1 Experimental details
For MinerU, we use version 0.9.2101010https://github.com/opendatalab/MinerU/releases/tag/magic_pdf-0.9.2-released by default. For Marker, version 0.2.17111111https://github.com/VikParuchuri/marker/releases/tag/v0.2.17 is employed. For Nougat, we utilize its 0.1.0-base model (350M). All prompt templates can be found in Sec. I.
@@
IV.2 Details in different domains
@@ 2. Formula Processing:
[…]
@@ V Case Study
Fig. S2 to  Fig. S11 show some cases of GOT, MinerU, and Qwen2.5VL-72B on OHRbench. For each case, we indicate the evidence source and answer, giving the OCR result of different models and the responses at the retrieval and generation stages.

### v3

@@ 4.3 In-depth analysis of OCR noise’s impact on RAG
In this section, we conduct an in-depth analysis of the impact of Semantic Noise and Formatting Noise on RAG systems, using perturbed structured data with varying levels of perturbations. For each type of OCR noise, we introduce three noise levels—mild, moderate, and severe—to systematically assess their effects. Since our findings in Sec.˜4.2 indicate that edit distance fails to accurately capture the degree of OCR noise, we instead define the ratio $r_{\text{noise}}$, representing the proportion of Q&A pairs affected by OCR noise as a measure. Specifically, for each Q&A pair, we compute the LCS between its evidence context and the corresponding perturbed structured data. If LCS exceeds 0.95, the Q&A pair is considered as unaffected, otherwise, it is affected. We then use the ratio ($r_{\text{noise}}$) to quantify the degree of perturbations, with 0 representing to ground truth structured data and values approaching 1 indicating greater perturbation. This approach allows us to align the perturbation levels of Formatting Noise and Semantic noise for fair comparisons. Additionally, for Formatting Noise, we evaluate its retrieval performance with modified LCS calculations by excluding stylistic commanding introduced during perturbation, ensuring a fair assessment of retrieval accuracy. For each degree of Semantic Noise, we report the average RAG performance using three different OCR results, including MinerU, GOT, and Qwen2.5-VL-72B.
@@ 4.3.1 Fine-grained impact on retrieval and generation
Semantic Noise significantly influences both retrieval and generation phases. As illustrated in Fig.˜3, increasing Semantic Noise from mild ($r_{\text{noise}}=0$) to severe ($r_{\text{noise}}>0.6$) results in nearly a $50\%$ performance decline for most retrievers and LLMs. In the retrieval stage, both the sparse retriever BM25 and the dense retriever BGE-M3 suffer consistent performance declines across all types of questions, suggesting that dense retrieval’s stronger comprehension does not provide robustness against Semantic Noise. In the generation phase, all LLMs struggle with Semantic Noise, among which performance on reading-order decreases the most. Interestingly, although the way we introduce Semantic Noise should primarily affect text recognition, questions related to multimodal elements (tables, formulas and charts) degrade even further, highlighting the challenges in parsing, understanding and reasoning over multimodal document data.
@@ 4.3.2 Impact on end-to-end evaluation
Semantic Noise consistently demonstrate a strong impact, while Formatting Noise affects specific retrievers and LLMs differently. Semantic Noise consistently degrades performance across all combination of retrievers and LLMs, particularly on multimodal questions involving tables, formulas, and charts. In contrast, the effect of Formatting Noise is more variable when using smaller LLMs, Llama3.1-8B and Qwen2-7B, for generation, despite greatly reduced retrieval accuracy, the overall performance shows a slight change due to their limited information integration capabilities. Conversely, using a larger LLM, Qwen2-72B, is highly sensitive to retrieval performance. While its generation performance shows light changes, the overall performance decreases, especially on questions related to multimodal elements.
In summary, Semantic Noise significantly affects each stage of RAG and the entire system. The impact of Formatting Noise varies with different retrievers and generators, particularly affecting questions with multimodal elements.
@@ 4.3.3 Impact of table format
In addition to perturbations, we investigate the influence of different table formats as a kind of Formatting Noise. As shown in Fig.˜4, HTML tables show inferior performance during retrieval compared to the Markdown and LaTeX formats. Markdown and LaTeX formats perform similarly, with BGE-M3 demonstrating a better understanding of Markdown. In the generation phase, HTML and LaTeX showed similar performance across all models, but the Markdown format performed worse due to the lack of support for merging cells. In end-to-end evaluations affected by low retrieval performance, using HTML tables is comparatively worse, while the combination of Qwen2-72B and BGE-M3 using Markdown achieves the best performance.
@@ 4.3.4 Error analysis
We further conduct error analysis to understand the bottlenecks of the RAG system with OCR noise in a quality approach. Specifically, we calculate the distribution of OCR errors and RAG errors when using ground truth structured data and perturbed data with severe Formatting Noise and Semantic Noise. Our evaluation employs BGE-M3 as the retriever and assesses error distributions using two generators: Qwen2-7B and Qwen2-72B. We use the same strategies to identify the proportion of OCR errors as in Sec.˜4.3. The distribution of these errors is illustrated in Fig.˜5. It indicates that when the proportion of OCR errors is similar for Semantic Noise and Formatting Noise (66% vs. 61%), it performs better with Formatting Noise. Of these, about half of the correct responses when using Qwen2-7B as a generator were done by the model with incorrect OCR results. Meanwhile, larger LLMs are more robust to OCR noise. Compared to Qwen2-7B, Qwen-72B has a 9.4% and 3.5% higher percentage of samples with OCR errors but ultimately correct in Formatting Noise and Semantic Noise, respectively.
@@ 5 Conclusion
Acknowledgments
This work is supported by Shanghai Artificial Intelligence Laboratory, the National Key R&D Program of China (2024YFA1014003), the National Natural Science Foundation of China (Grants 92470121 and 62402016), and the CAAI-Ant Group Research Fund.
@@ I Instruction Prompts
Q&A Generation Prompt Template. The template is shown in Tab.˜S9. Following [53, 58], we instruct GPT-4o to generate questions with clear entities and require three levels of difficulty for question diversity.
RAG Generation Stage Prompt Template. The prompt template for LLMs and VLMs with text-only input is shown in Tab.˜S12.
Vision-Language Models OCR Prompt Template. We tune the prompt for the best performance of VLMs OCR, by comparing simple and detailed instructions as shown in Tab.˜S13. Results in Tab.˜S14 indicate that the detailed prompt consistently performs better across all evaluations, so it is used by default.
@@ II.1 Document details
DUDE: We extract documents from the validation and test splits of DUDE, applying manual screening based on our criteria Fig.˜1 to exclude samples infeasible for structured data parsing and classify each of them into 7 domains. We finally selected 450 PDFs with 4,058 images from 2,069 PDF candidates.
@@
Each document is manually reviewed by primary authors to ensure its availability for academic use. Detailed domain statistics are shown in Tab.˜S1
@@ II.4 Q&A generation
To generate high-quality Q&A pairs covering diverse tasks and evidence sources, we define multiple prompts for each task, as detailed Tabs.˜S9, S10 and S11. For Chinese questions, we provide the same set of templates in Chinese to ensure that the model generates Chinese responses. Q&A with different evidence sources. For Q&A generation with evidence sourced from plaint text, table, formula and chart, we extract relevant pages from the ground truth structured data and use GPT-4o to generate Q&A pairs grounded in the corresponding evidence via tailored prompts. For Q&A related to reading order, we leverage MinerU [48], the leading model for reading order recognition [35], to identify the reading order and bounding box of paragraphs in each document. When working with documents from OmniDocBench [35], we directly use the ground truth reading order from its annotations. We verify the layout detection and reading order predictions, selecting paragraph pairs that meet one of the following criteria:
@@
We then randomly sample 1,500 candidate matches, manually correcting approximately 20% where MinerU’s predictions are inaccurate. We then prompt GPT-4o to generate Q&A pairs using the prompts in Tab.˜S10. We find that this simple prompting-based strategy can effectively generate questions with diverse evidence sources, with over 90% correctly aligned with their evidence source in our Q&A verification process.
Q&A with different tasks. To generate both understanding and reasoning questions, we apply the corresponding prompts from Tab.˜S10. For multi-page Q&A generation, we employ two different approaches to generate Q&A candidates: (1) Combine questions from two single-page Q&As that mention the same entity. (2) Generating multi-page questions from two paragraphs on different pages that reference the same entity. Specifically, we use spaCy [19] for named entity recognition in both single-page Q&As and document paragraphs. We then filter out candidate pairs, including: (1) Single-page Q&A pairs where the entity in one answer appears in another question. (2) Paragraph pairs that share the same entity. We finally utilize the prompts in Tab.˜S11 to generate multi-page questions. However, despite the many optimizations of the prompt and generation strategies we tried, GPT-4o sometimes produces Q&A pairs that are either answerable with a single paragraph or simply concatenate two single-page questions while maintaining separate evidence sources instead of high-quality and realistic multi-page Q&As. To address these limitations, we develop a comprehensive filtering process to ensure the quality of multi-page Q&As, as detailed in Sec.˜II.5.
@@
Faithfulness to Task Definition. Based on the Q&A verification prompts in [12], we use the prompts in Tab.˜S15 to assess faithfulness using DeepSeek-V3. To verify the validity of evidence sources, we locate them in the original ground truth structured data and ensure they originate from the correct corresponding LaTeX code environments. For the multi-page and reading-order questions, we employ GPT-4o to generate three responses: (1) without context, (2) with context A, and (3) with context B. If any response yields a correct answer, the question is excluded, ensuring that only truly multi-page or reading-order-related questions remain.
@@ III.2 Rules for Semantic Noise introduction
  • Random Rotation: Apply a random rotation of $3^{\circ}$ and $3^{\circ}$.
@@
We generate three document image datasets with varying noise levels and parse structured data using MinerU, GOT, and Qwen2.5-VL, resulting in 9 perturbed datasets. The examples of distorted documents are shown in Fig.˜S1. The distribution of introduced Semantic Noise is illustrated in Fig.˜S2. In most cases, the distributions of our perturbed PDFs align with those of real-world distorted PDFs, validating the realism of our method. In Sec.˜4.3, we evaluate RAG performance on these datasets, reporting the average results for each noise level.
Ratio of OCR noise in real-world OCR results. To illustrate the frequency of OCR noise in real-world OCR results, We match corresponding TXT/FOR/TAB blocks, which includes ~130 tokens each, and show the ratio of Semantic Noise and Formatting Noise in Tab.˜S3.
OCR SN FN
TXT FOR TAB
MinerU 40.3% 78%/34% 79%/58% 10.9
Qwen2.5-VL 31.6% 46%/25% 75%/60% 16.4
Table S3: SN: ratio of matching blocks with textual/structural errors. FN: average redundant formatting commands per page.
Figure S1: Cases of distorted documents.
Figure S2: Distribution of Semantic Noise. X-axis denotes edit distance. Mild/moderate/severe perturbation is based on born-digital PDFs.
@@ IV.1 Experimental details
For MinerU, we use version 0.9.2101010https://github.com/opendatalab/MinerU/releases/tag/magic_pdf-0.9.2-released by default. For Marker, version 0.2.17111111https://github.com/VikParuchuri/marker/releases/tag/v0.2.17 is employed. For Nougat, we utilize its 0.1.0-base model (350M). All prompt templates can be found in Sec.˜I.
@@
IV.2 Sim2Real GAP
As the questions posed by human users could have far more diversity in styles than LLM generated Q&As. We randomly pick 100 Q&As and manually rewrite questions for comparison. The performance before and after rewriting is: 27.2/23.2(GT), 20.7/18.0(MinerU), 12.8/12.9(GOT), and 23.1/20.0(Qwen2.5-VL). Although performance degrade, the conclusions about different OCR solutions still hold, as question styles may primarily be associated with models’ ability to understand instructions.
IV.3 Multimodal RAG
We compare VisRAG with OCR-based RAG, using Qwen2-VL-7B as the generator for fair comparison. The results are shown in Tab.˜S4. VisRAG achieves competitive results on multimodal element-related Q&As (e.g. table and chart), but underperforms on TXT and RO (e.g. high-resolution newspapers), exhibiting similar failure modes to Qwen2.5-VL.
[…]
GT 57.7 41.7 41.8 39.8 29.8 46.8
MinerU 51.2 33.7 32.1 10.7 29.8 37.9
Qwen2.5-VL 54.1 38.9 34.5 22.7 13.3 40.6
VisRAG 50.7 40.3 32.0 30.6 15.6 40.2
Table S4: Performance of VisRAG and OCR-based RAG. We use Qwen2-VL-7B as the generator for fair comparison.
IV.4 Effectiveness of robust generator
We employ Ext2Gen-8B-R2 [41] and show its performance in Tab.˜S5. Ext2Gen-8B-R2 consistently improves performance. Although it is based on Llama3.1-8B, its performance on Azure remains stable, reinforcing that stronger models exhibit greater robustness to formatting noise. This further supports our conclusion that stronger models are more robust to formatting noise. However, the performance gap between the best OCR (Azure) and GT also increases by 1.6 compared to Qwen2-7B, indicating that OCR quality becomes a bottleneck and leaves a room for improvement.
[…]
Generator: Qwen2-7B/Llama3.1-8B
GT - 46.7/43.1 31.8/37.4 27.6/28.4 31.1/34.7 23.7/13.7 36.2/35.9 -
MinerU 0.24 42.2/37.8 27.0/30.0 23.5/22.5 8.9/9.7 23.0/12.5 30.5/28.5 -5.7/-7.4
Qwen2.5-VL 0.18 42.5/38.6 29.1/33.1 26.1/26.1 18.5/19.6 10.9/6.7 31.5/30.7 -4.7/-5.2
Azure 0.17 45.5/29.6 30.7/25.4 23.3/21.9 19.1/11.0 23.5/11.5 33.8/24.0 -2.4/-11.9
Generator: Ext2Gen-8B-R2
GT - 56.3 45.4 40.7 38.9 27.4 46.8 -
MinerU 0.24 49.7 36.4 30.5 10.8 25.8 37.6 -9.2
Qwen2.5-VL 0.18 52.7 41.2 34.8 24.6 12.9 40.9 -5.9
Azure 0.17 55.1 41.9 32.4 23.4 26.0 42.8 -4.0
Table S5: Experiments of Azure and Ext2Gen.
IV.5 Commercial OCR
We evaluate Azure OCR in Tab.˜S5 and observe the following: With powerful generators (Qwen2-7B and Ext2Gen-8b-R2), Azure yields the best performance, though there remains a gap of up to 4.0 compared to GT. But, when using Llama3.1-8B, performance drops significantly, even worse than MinerU. Our manual check suggests this may be due to custom formatting tags in Azure’s outputs, affecting Llama3.1-8B’s generation.
IV.6 Details in different domains
Tab.˜S6Tab.˜S7 and  Tab.˜S8 shows the performance of different OCR solution on different domains respectively.
@@ 2. Formula Processing:
 - Enclose block formulas with \$\$ \$\$. For example: \$\$ $\frac${-b $\pm$ $\sqrt${b2̂ - 4ac}}{2a} \$\$.
@@ V Case Study
Fig.˜S3 to  Fig.˜S12 show some cases of GOT, MinerU, and Qwen2.5VL-72B on OHRbench. For each case, we indicate the evidence source and answer, giving the OCR result of different models and the responses at the retrieval and generation stages.